The table below represents the hourly output and cost structure for Noah and Matthew’s pizza shop which operates in a perfectly competitive market. The market price of a pizza is $10.

1. Complete the table
2. Since the market price is $10, what is their short run profit?
3. What is their break even price? Explain
4. Suppose the market price decreases to $5 per pizza. In the short run, should Noah and Matthew continue to operate the pizza shop, or will they maximize their economic profits (that is, minimize their economic loss) by shutting down? Explain

Total Cost ($) AFC AVC ATC MC Output per hour . UN U
Answer #1

AFC=5/Q

AVC=VC/Q

ATC=TC/Q

MC= change in TC/change in Q

Q TC AFC AVC ATC MC
0 5
1 9 5.00 4.00 9.00 4
2 11 2.50 3.00 5.50 2
3 12 1.67 2.33 4.00 1
4 14 1.25 2.25 3.50 2
5 18 1.00 2.60 3.60 4
6 24 0.83 3.17 4.00 6
7 32 0.71 3.86 4.57 8
8 42 0.63 4.63 5.25 10
9 54 0.56 5.44 6.00 12
10 68 0.50 6.30 6.80 14

2) P=10

Setting P=MC the firm will produce 8 units so profits = TR-TC = 80-42 = 38

3) At breakeven total revenue = total cost which would occur when the Price = minimum ATC = 3.50

4) In the short run at P=5, the firm should continue to produce as price is greater than minimum AVC

They are able to cover their variable costs so they should keep producing and only shut down in the long run